8: A: An electron and a proton are released from rest in a constant electric field with a magnitude of 200N/w. Compare the magnitude and direction of the force felt by each charge. Ignore the small attractive force between the electron and proton. B: Compare the magnitude of the acceleration of each charge. How far does each charge move in the first 1.0us ? Use m,-1.67x10 kg and m,-9.11x10kg
Answer #1

8) A) force on proton = qE = 1.6*10^-19*200

= 3.2*10^-17 N (in the right direction)

force on proton = qE = 1.6*10^-19*200

= 3.2*10^-17 N (in the left direction)

B) acceleration of proton = F/m = 3.2*10^-17 / 1.67*10^-27 = 1.92*10^10 m/s2

S= ut + 0.5at^2

S = 0 + 0.5*1.92*10^10*(10^-6)^2

= 9.5 mm

acceleration of electron = F/m = 3.2*10^-17 / 9.11*10^-31

= 3.5*10^13 m/s2

S= ut + 0.5at^2

S = 0 + 0.5*3.5*10^13*(10^-6)^2

= 17.56 m